OpenStudy (anonymous):

a right triangle has a perimeter of 3m. its hypotenuse is 130cm. ehat are the lengths of the other two sides.

OpenStudy (anonymous):

perimeter is 300 cm 300 - 130 = 170 a^2 + b^2 = 16900 a + b = 170 a = 170 - b (28900 - 340b + b^2) + b^2 = 16900 2b^2 - 340b + 12000 = 0 b^2 - 170b + 6000 = 0 (b - 50)(b-120) b = 50, 120 The two sides are 50 and 120.